Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by recocollins » Tue Jul 31, 2007 10:29 am
(1) In a drawer of shirts 8 are blue, 6 are green, and 4 are magenta. If Mason draws 2 shirts at random, what is the probabililty at least one of the shirts he draws will be blue?

A. 25/153
B. 28/153
C. 5/17
D. 4/9
E. 12/17

(2) A certain consulting firm employs 8 men and 4 women. In March, 3 employees are selected at random to represent the company at a convention, what is the probability that the representatives will NOT all be men?

A. 14/55
B. 3/8
C. 41/55
D. 2/3
E. 54/55

(3) Kurt, a painter, has 9 jars of paint 4 of which are yellow, 2 are red and the remainng jars are brown. Kurt will combine 3 juars of paint into a new container to make a new color which he will name according to the following conditions:

Brun Y if the paint contains 2 jars of brown paint and no yellow
Brun X if the paint contains 3 jars of brown paint
Jaune X if the paint contains at last 2 jars of yellow
Jaune Y if the paint contains exactly 1 jar of yellow

What it the probability that the new color will be Juane?

A. 5/42
B. 37/42
C. 1/21
D. 4/9
E. 5/9


I greatly appreciate the help around these questions. :)
Join the discussion
Source: — Problem Solving |

by RM » Tue Jul 31, 2007 10:58 am
1. In a drawer of shirts 8 are blue, 6 are green, and 4 are magenta. If Mason draws 2 shirts at random, what is the probabililty at least one of the shirts he draws will be blue?

Blue 8; Green 6; Magenta 4; Total 18

Total Number of Ways to draw 2 shirts = 18C2
Number of Ways to draw NO Blue shirts = (6+4)C2 = 10C2

Probability of drawing NO Blue Shirt = 10C2/18C2 = 45/153 = 5/17
Probabilty of drawing atleast 1 Blue Shirt = 1 - 5/17 = 12/17
Choice E

2. A certain consulting firm employs 8 men and 4 women. In March, 3 employees are selected at random to represent the company at a convention, what is the probability that the representatives will NOT all be men?

"Probability that the representatives will NOT all be men" is same as "Probability that the representatives will have atleaset 1 women"

Men 8; Women 4; Total 12

Total Number of ways to select 3 employees from group = 12C3
Number of ways to select ALL men = 8C3

Probablity of selecting all Men = 8C3/12C3 = 14/55
Probability of selecting NOT all Men Or atleast One Women = 1 - 14/55 = 41/55. Choice C.
Join the discussion

by recocollins » Tue Jul 31, 2007 12:00 pm
Thanks a ton. It all makes more sense now.
Join the discussion

by recocollins » Tue Jul 31, 2007 12:10 pm
Thanks a ton. It all makes more sense now.
Join the discussion

by jay2007 » Tue Jul 31, 2007 12:49 pm
For 1 and 2 RM is correct:

For 3:
=====
(3) Kurt, a painter, has 9 jars of paint 4 of which are yellow, 2 are red and the remainng jars are brown. Kurt will combine 3 juars of paint into a new container to make a new color which he will name according to the following conditions:

Brun Y if the paint contains 2 jars of brown paint and no yellow
Brun X if the paint contains 3 jars of brown paint
Jaune X if the paint contains at last 2 jars of yellow
Jaune Y if the paint contains exactly 1 jar of yellow

What it the probability that the new color will be Juane?

Juane = J say.

Probability of J = Jx + Jy

Jy (probability of 1 exactly jar of yellow) = ((4c1)*(5c2))/(9c3) = 10/21
Jx = (probability of exactly 2 jars of yellow) + probability of exactly 3 jars of yellow
= ((4c2)*(5c1))/(9c3) + ((4c3)*(5c0))/(9c3)
= 17/42
therefore j = 10/21 + 17/42 = 37/42

Therfore answer is B.
Join the discussion